Find the value of "r" so the line that passes through each pair of points has the given slope.
(4,-5), (3,r) m=8
My answer is -3. Am I correct?
Here is how I got my answer: r - -5 over 3 - 4
8 = r- -5
_________
1 -1
8(-1)= 1(r - -5)
-8= 1r - -5
+5 +5
____________
-3=r

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

Find the value of "r" so the line that passes through each paur of points has the given slope.
(4,-5), (3,r) m=8
My answer is -3. Am I correct?
-----------------------
(r--5)/(3-4) = 8
(r+5)/-1 = 8
r+5 = -8
r = -13
============
Cheers,
Stan H.

Answer by edjones(8007)

point-slope form y-y[1]=m(x-x[1])
m=8 let r=y, -5=y[1], 3=x, 4=x[1] you can do it the other way and the answer will be the same.
r-(-5)=8(3-4)
r+5=-8
add -5 to both sides
r= -13
.
Ed